# Cosmological Roles of Dark Photons in Axion-induced Electroweak Baryogenesis

### v2

@@ IV Conclusions
If the SM is extended to include an axion that couples to both the Higgs and electroweak gauge sectors, the matter-antimatter asymmetry in the universe can be explained through electroweak baryogenesis. For this to occur, the axion decay constant should lie between approximately $10^{5}$ and $10^{7}$ GeV, corresponding to axion masses ranging from the MeV to GeV scales. It is plausible that the axion responsible for baryogenesis interacts with dark photons via a natural coupling arising from the U$(1)$ anomaly. These axion-coupled dark photons can suppress the branching ratio of axion decay to SM particles, thereby relaxing the constraints on the axion that arise from supernovae explosions and BBN. Moreover, axion-coupled dark photons can constitute the main component of dark matter as they are produced from the decay of axions in thermal bath through the freeze-in mechanism. Additionally, ultralight axion-coupled dark photons may provide an observable contribution to dark radiation. Unlike other scenarios for electroweak baryogenesis, our model can be probed through searches for axion-like particles rather than relying on EDM and LHC experiments. This is because the axion interacts with the SM feebly, via couplings suppressed by a large decay constant.

### v3

@@ IV Conclusions
If the SM is extended to include an axion that couples to both the Higgs and electroweak gauge sectors, the matter-antimatter asymmetry in the universe can be explained through electroweak baryogenesis. For this to occur, the axion decay constant should lie between approximately $10^{5}$ and $10^{7}$ GeV, corresponding to axion masses ranging from the MeV to GeV scales. It is plausible that the axion responsible for baryogenesis interacts with dark photons via a natural coupling arising from the U$(1)$ anomaly. These axion-coupled dark photons can suppress the branching ratio of axion decay to SM particles, thereby relaxing the constraints on the axion that arise from SN explosions and BBN. Moreover, axion-coupled dark photons can constitute the main component of dark matter as they are produced from the decay of axions in thermal bath through the freeze-in mechanism. Additionally, ultralight axion-coupled dark photons may provide an observable contribution to dark radiation. Unlike other scenarios for electroweak baryogenesis, our model can be probed through searches for axion-like particles rather than relying on EDM and LHC experiments. This is because the axion interacts with the SM feebly, via couplings suppressed by a large decay constant.